### v1

@@ 4.1 Single-stage Approaches
  • For Sign Language Understanding, the authors in [157] developed G-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in [54] proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2D-CNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER[93] to adopt it for ISL recognition, while authors in [161] leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP[8] has also been developed, to simultaneous detect and identify the hand gesture.
@@
  • For gesture-based Human-Computer Interaction the authors in [198] proposed a lightweight 2D-CNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in [80] proposed 2DPSTPP-Net, a 2D-CNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in [61] modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes.
  • General-purpose models have also been developed. EDenseNet [171] extends DenseNet by adding convolutional layers between dense blocks. SpAtNet [21] is a lightweight 2DCNN with three streams: two MAFF modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. MMFANet [103] performs gesture detection and recognition using a 2DCNN backbone with task-specific regressor and classifier branches.
4.2 Multi-stage Approaches
Figure 8: Architecture of an indicative multi-stage dual-stream approach for static VHGR named FGDSNet [208].
In contrast to single-stage, multi-stage approaches set up a multi-model processing pipeline. Typically, a hand detector or segmentation model is used to remove the background and normalize the spatial dimensions. These methods are primarily applied in Human Robot Interaction (HRI) and HCI scenarios, as detailed below:
@@ 4.2 Multi-stage Approaches
  • Human-Robot Interaction, particularly gesture-based mobile robot navigation, has attracted research interest. URGR [17] is a three-stage approach using YOLOv3 for human cropping, HQ-Net for super-resolution to address long-range scenarios, and a classifier for gesture recognition. Similarly, [153] introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections.
  • Multi-stage static VHGR methods have also been proposed for Human-Computer Interaction. In [9], a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In [34], a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2D-CNN like MobileNet or EfficientNet. [208] introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
4.3 State-of-the-art Methods for Static VHGR
The current SOTA static VHGR methods are summarized in 7, based on input modality, architecture, and performance on public datasets, including HGR1 [88], NUS-I [95], NUS-II [136], ASL-FS [137], Jochen-Triesch II [175], and OUHANDS [118] [^1] For datasets with splits for both subject-independent and subject-dependent modes, only the latter is considered.. Since most methods are evaluated on different or custom datasets, a straightforward performance comparison is not possible.
As shown, most SOTA methods rely on RGB data, with depth-based DRCAM [153] lagging behind RGB methods on ASL-FS. Researchers use attention mechanisms [153, 198, 208], multi-stream approaches [21, 208], and segmentation modules for background removal [208, 34, 153] to improve performance. On OUHANDS [118], ASL-FS [137], and NUS-I [95], single-stage methods outperform multi-stage ones, while on NUS-II [136], FGDSNet [208] with segmentation achieves slightly better accuracy than 2DPSTPP-Net [80]. Thus, background removal does not always improve sign recognition, as single-stage methods often surpass multi-stage models.
Table 3: Static HGR methods and their performance on well-known datasets. The highest accuracies are shown with bold. Methods marked with star (*) perform hand segmentation.
Method Modality Architecture Description Dataset Accuracy
SpAtNet [21] RGB Three-stream 2DCNN, where each stream aims to obtain either high-level or coarse-grained features HGR1 93.33%
ASL-FS 99.67%
NUS-II 96.53%
FGDSNet [208]* RGB Dual stream 2DCNN approach and a custom dual-stream segmentation module to obtain the mask HGR1 94.8%
OUHANDS 97.8%
NUS-II 99.8%
[34]* RGB DL-based Hand Segmenation + Classifier HGR1 98.15%
OUHANDS 89.2%
LHGR-Net [198] RGB BaseNet + MSS (adopting the atrous spatial pyramid pooling) + LAS (based on CBAM) HGR1 93.15%
OUHANDS 98.75%
2DPSTPP-Net [80] RGB Convolutional blocks followed by the proposed pooling method, intermediate features are concatenated to obtain the final representation. NUS-I 100%
NUS-II 99.61%
DRCAM [153]* Depth Attention-based 2CNN with residual connections ASL-FS 93.44%
Gesture-CNN [157] RGB 2DCNN consisting of four convolutional layers Jochen-Triesch II 100%
EDenseNet [171] RGB DenseNet with extra convolution layers ASL-FS 99.99%
NUS-II 99.3%
@@ 5 Methods for Isolated Dynamic VHGR
Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
@@
In the following, we present IDVHGR methods grouped by input modality as (a) unimodal appearance-based, (b) pose-based, and (c) multimodal approaches. Then, we conclude with a summary of key innovations and state-of-the-art (SOTA) performance.
@@ 5.1.1 Spatiotemporal
In [182], the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in [20] proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3D CNN for classification.
In [186], a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in [63] introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, [173] addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
@@ 5.1.2 Spatial + Temporal
In [65], a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in [178] introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in [183]. The model uses a 2D-CNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in [40] is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
@@ 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
EgoFormer [138] is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3D-CNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in [18] proposed a pipeline where a lightweight 2D CNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, [22] proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
Figure 9: The zero-shot framework proposed by authors in [22]. Each branch of the dual-stream backbone consists of a spatial-temporal model, such as the I3D, followed by a long-term recurrent model. The BERT is utilized to map textual description into the embedding space. These features are also combined with hand-crafted attributes.
@@ 5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in [113] for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2D-CNN. A different approach was presented in [94], where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, [110] proposed a dual-stream model with keyframe selection. A 2D-CNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer [109], another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
5.2 Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination [38]. They also explore joint subsets related to specific body regions [135], often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition [125]. Consequently, many models integrate additional cues—typically from RGB inputs—using 3D CNNs and perform score-level fusion to enhance performance.
@@ 5.2.1 Spatiotemporal
Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU), sign language recognition (SLR), and human-computer interaction (HCI).
In Sign Language Understanding (SLU), pose-based methods aim to capture dynamic hand and body movements for gesture recognition. SKIM [104] uses a dynamic graph convolutional network (GCN) combined with temporal convolution (TCN) to analyze both global and local body regions, with a feature extraction module focusing on high-motion joints. Similarly, MC-LSTM [133] employs multiple streams (face, hand, and whole body) processed by temporal convolutions and GCN for improved feature ensembling. Another notable method, P3D [98], uses a transformer-based model with part-wise and whole-body encoders to process landmark data for classification. For improved performance, HWGAT [135] utilizes hierarchical spatial windows with attention mechanisms to highlight key body parts. TMS-Net [38] proposes a six-stream GCN-based approach for processing skeleton joints, bones, and motion data, with inter-stream information exchange. Additionally, methods like SLGTformer [168] and VSCT [25] address the challenge of data scarcity by leveraging transformers and few-shot learning, respectively.
In Sign Language Recognition (SLR), pose-based methods often focus on improving the accuracy and robustness of sign classification. Authors in [161] present a dual-stream GCN model for Korean SLR, analyzing joint and motion data. Similarly, MULTISTAGE-CGAR [121] proposes a dual-stream architecture with separable temporal convolution and attention mechanisms for improved feature extraction. For multi-stream fusion, SwC GR-MMixer [109] combines GCN with GRU units for temporal modeling and attention mechanisms to reduce overfitting. Notably, L-TSSI-DenseNet [97] explores skeleton graph representations by converting joint coordinates into image-like matrices processed by DenseNet for classification. Moreover, STST [201] leverages a two-stage transformer-based approach for global and local spatial-temporal feature extraction, offering a dynamic alternative to traditional methods.
In Human-Computer Interaction (HCI), pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks [79] use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN [167] proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA [5], tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN [30] addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Additionally, DSTA-SLR [77] improves spatiotemporal modeling by introducing dynamic adjacency matrices and multi-stream fusion, enhancing the accuracy of joint, bone, and motion-based recognition tasks.

### v2

@@ 4.1 Sign Language Understanding (SLU)
For SLU, the authors in (G-CNN-SHARMA2021115657) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in (STN-ND-Net) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER (SIGNFORMER) to adopt it for ISL recognition, while authors in (Jungpil-Shin) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP(YOLOv4-CSP) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
4.2 Human-Computer Interaction (HCI)
@@
For gesture-based HCI the authors in (LHGR-Net) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in (2DPSTPP) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in (WEI_HAN_YOLOV5_COPULA) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes. In (Mohammad_Mahmudul_static_hgr), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In (tuan_linh_pham_dang2023lightweight), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet. (FGDSNet) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
4.3 Human-Robot Interaction (HRI)
@@
HRI, particularly gesture-based mobile robot navigation, has attracted research interest. URGR (ULTRA-RANGE-UGV-Bamani) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition. The HQ-Net comprises convolutional and deconvolutional pathways. Features extracted by Canny Edge algorithm and a convolutional module followed by self-attention are fused with the latent representation. GViT utilizes two GCN layers that capture fine-grained details followed by a modified ViT. This setting enables effective recognition of gestures in a distance of up to 25m for UGV guidance. Similarly, (DRCAM_TIM) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks. Because of the attention mechanisms, the model focuses effectively on the most informative regions of the image.
4.4 General Purpose
General-purpose models have also been developed. EDenseNet (EDenseNet) extends DenseNet by adding convolutional layers between dense blocks. SpAtNet (SpAtNet) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. This approach, employing multiple kernel sizes, increases accuracy by capturing various levels of granularity. MMFANet (MMFANet) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches. The FAPN module aims to extract multiscale features through upsampling/downsampling operations and fusion of adjacent feature maps, while reducing the computational complexity by utilizing novel expanded convolutional blocks with different dilation rates combined with attention mechanisms that eliminate redundant information.
4.5 Performance of SHGR Methods
Most existing methods for Static Hand Gesture Recognition (SHGR) are evaluated on different or custom datasets, which prevents a direct comparison of their performance. Table 2 presents the accuracy of current state-of-the-art SHGR approaches. Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (ASL-FS-2011-CVPRW), HGR1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands)) remain relatively small and lack diversity. Consequently, model performance would likely degrade in real-world, uncontrolled environments. Among the existing methods, LHGR-Net (LHGR-Net) and FGDSNet (FGDSNet) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction. On the other hand, EDenseNet (EDenseNet) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition. Finally, applying background removal through hand segmentation does not appear to significantly improve accuracy compared to end-to-end learning approaches.
Table 2: Performance of the most accurate SHGR methods. Many approaches achieve high accuracy even with small model size, but the limited size and diversity of the utilized datasets may result to performance reduction during real-world deployment. Methods marked with star (*) perform hand palm segmentation first.
Method Year Modality Params ASL-FS HGR1 NUS-I NUS-II OUHANDS
FGDSNet (FGDSNet) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
SpAtNet (SpAtNet) 2024 RGB 3.8M - 93.33% - - -
(tuan_linh_pham_dang2023lightweight) * 2023 RGB 5.34M - 98.15% - - 89.20%
LHGR-Net (LHGR-Net) 2023 RGB 1.81M - 93.15% - - 98.57%
2DPSTPP-Net (2DPSTPP) 2023 RGB 19.01M - - 100% 99.61% -
DRCAM (DRCAM_TIM) * 2023 Depth - 93.44% - - - -
EDenseNet (EDenseNet) 2021 RGB 7.77M 99.99% - - 99.30% -
@@ 5 Methods for Isolated Dynamic VHGR
Unlike static VHGR approaches, which often fail to capture temporal dynamics and struggle with motion-based gestures, recent advances in Isolated Dynamic Vision-based Hand Gesture Recognition (IDVHGR or IDHGR) have focused on recognizing individual gesture instances from spatiotemporal inputs such as RGB video sequences or skeleton joint trajectories.
@@
In the following, we present IDVHGR methods grouped by input modality as (a) unimodal appearance-based, (b) pose-based, and (c) multimodal approaches. Then, we conclude with a summary of key innovations and SoTA performance.
@@ 5.1.1 Spatiotemporal
In (2plus1-D-SLR_and_NCSL), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in (Snehal_Bharti_et_al) proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3DCNN for classification.
In (ShiqiWang_et_al), a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in (TS3C-Net) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, (SeST) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
@@ 5.1.2 Spatial + Temporal
In (DAVID_RICHARD_TOM_HAX), a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in (GoogleNet-BiLSTM) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (TIM-SLR). The model uses a 2DCNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in (Full_Transformer_Network) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
@@ 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
EgoFormer (EgoFormer_METHOD_EgoDriving_DATASET) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3DCNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in (SlowFast-Transformer) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, (ZERO-SHOT-SLR) proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
@@ 5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in (HAMZAH_LUQMAN) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2DCNN. A different approach was presented in (Ensemble-InceptionV4), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, (Tianyu_Liu) proposed a dual-stream model with keyframe selection. A 2DCNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer (SwC-GR-MMixer), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.
5.2 Unimodal Pose-based Methods
Pose-based methods, also known as skeleton-based approaches, rely on structured motion data—such as skeleton graphs and pose heatmaps—to classify gestures. These methods often incorporate various forms of pose information, including joint coordinates, bones, motion trajectories, and joint angles, sometimes in combination (TMS-NET). They also explore joint subsets related to specific body regions (FDMSE-hierarchicalwindowedgraphattention), often adopting multi-stream architectures to handle these diverse inputs.
Given the natural graph structure of skeletal data, many methods employ Graph Convolutional Networks (GCNs) originally developed for human activity recognition. In the sections that follow, we group the approaches based on the structure of their backbone, following the same organization as in the previous subsection. Notably, pose information alone is often insufficient for accurate sign recognition (moryossef2021evaluating). Consequently, many models integrate additional cues—typically from RGB inputs—using 3DCNNs and perform score-level fusion to enhance performance.
@@ 5.2.1 Spatiotemporal
Pose-based methods that incorporate spatiotemporal representations simultaneously process spatial and temporal data to improve sign language understanding (SLU) and human-computer interaction (HCI).
In SLU, pose-based methods model dynamic hand and body movements for gesture recognition. SKIM (SKIM) combines dynamic graph convolutional (GCNs) and temporal convolutional (TCNs) networks to analyze global and local motion, while MC-LSTM (ozdemir2023multi_MC-LSTM) fuses multiple streams (face, hand, body) via temporal convolutions and GCNs. P3D (P3D_ICCV) employs transformer-based part-wise and whole-body encoders, and HWGAT (FDMSE-hierarchicalwindowedgraphattention) enhances key body-part focus with hierarchical attention. TMS-Net (TMS-NET) introduces a six-stream GCN for skeleton, bone, and motion data with inter-stream fusion. To tackle data limitations, SLGTformer (SLGTFORMER) and VSCT (spoter_vsct) use transformers and few-shot learning, respectively. Dual-stream approaches such as (Jungpil-Shin) and MULTISTAGE-CGAR (MULTISTAGE-CGAR) integrate joint and motion data through temporal convolutions and attention. SwC GR-MMixer (SwC-GR-MMixer) combines GCNs with GRUs for temporal modeling and regularization. Other notable advances include L-TSSI-DenseNet (L-TSSI-DenseNet), which converts joint coordinates into image-like matrices for DenseNet processing; STST (STST), a two-stage transformer for local–global spatiotemporal features; and DSTA-SLR (DSTA_SLR_hu2024dynamic), which employs dynamic adjacency matrices and multi-stream fusion for improved recognition accuracy.
In HCI, pose-based methods address challenges such as data overfitting and the need for self-supervised learning frameworks. MAE-based frameworks (omar_ikne) use a masked autoencoder to reconstruct skeleton graphs, focusing on motion and spatial relationships. 3sISTGCN (JOINTS_MOTION_VECTORS_Improved_ST-GCN) proposes a three-stream model to process joints, bones, and motions, applying modified ST-GCN layers and late fusion for improved classification. Another approach, DG-STA (BOAT-MI_VR), tackles class-incremental learning in HCI by training models on non-overlapping tasks without access to old training data. DSTSA-GCN (DSTSA-CGN_cui2025dstsa) addresses temporal expression limitations by stacking three stages of graph convolutions and multiscale temporal convolutions. Finally, authors in (ST-CGNet-hu2025st) proposed ST-CGNet, a dual-stream architecture comprising a 3DCNN and a convolutional LSTM module to capture different levels of information. Attention mechanisms and an advanced fusion strategy are utilized to efficiently incorporate the features of each stream.